A rectangle has length 36 cm and width 28 cm. Find the length of another rectangle of equal area whose width is 21 cm.
Answer by mananth(16949)

the length and width of the rectangle are 36 & 28 cm.
Area = L*W
Area = 36*28 = 1008 sq. cm
..
the width of other rectangle = 21cm
area = 1008
so length = area / width
length = 1008/21
= 48 cm
length = 48 cm